OpenStudy (anonymous):

Elena's daughter sold 84 boxes of girl scout cookies. She was told that this amount represented 5% of the troop's sale. How many boxes did the troop sell

OpenStudy (diyadiya):

Let the total number of boxes sold by the troop be \(\Large x \) 5% = \( \Large \frac{5}{100} \) \[\Large \frac{5}{100} \times x = 84\] Now solve for x

OpenStudy (diyadiya):

Can you try?

OpenStudy (anonymous):

168,000 is that right

OpenStudy (diyadiya):

Nope, But you're close can you show me how you got that?

OpenStudy (anonymous):

Would it be 1680

OpenStudy (diyadiya):

Yes, That's Right! :)

OpenStudy (anonymous):

Thank u I got another one jason is a sale person in the electronics department of a store. He is 48% short of his goal for the week. His goal is to earn $375. For each cd player he sells, he earns $45. How many cd players does he need to sell in order to meet his goal can u show me how to do it

OpenStudy (lgbasallote):

@Corya don't forget to medal the person who has given you the best answer which satisfied you..it is very customary here as a sign of respect and appreciation ^_^

OpenStudy (diyadiya):

lol :D

OpenStudy (anonymous):